# Rho-1: Not All Tokens Are What You Need

### v3

@@ 1 Introduction
Based on these analyses, we introduce Rho-1 models trained with a novel Selective Language Modeling (SLM) objective. As shown in Figure 2 (Right), this approach inputs the full sequence into the model and selectively removes the loss of undesired tokens. The detailed pipeline is depicted in Figure 4: First, SLM trains a reference language model on high-quality corpora. This model establishes utility metrics to score tokens according to the desired distribution, naturally filtering out unclean and irrelevant tokens. Second, SLM uses the reference model to score each token in a corpus using its loss (§2.2). Finally, we train a language model only on those tokens that exhibit a high excess loss between the reference and the training model, selectively learning the tokens that best benefit downstream applications (§2.2).
@@ B.2 Data Selection
Sample-level selection has been extensively studied in previous research, particularly through online batch selection methods (Loshchilov and Hutter, 2015; Schaul et al., 2015; Katharopoulos and Fleuret, 2018; Jiang et al., 2019). These approaches have been successfully applied to diverse classification tasks (Mindermann et al., 2022) and language modeling (Fan and Jaggi, 2023).
Token-level training strategies have also been explored, especially for the pre-training of BERT-like models using Masked Language Modeling (MLM) (Devlin et al., 2018). Specifically, “selective masking” involves masking important tokens in the input to focus on learning tokens that are more relevant to downstream tasks (Gu et al., 2020; Lad et al., 2022), whereas “token dropping” aims to reduce training costs by omitting less important tokens (Zhong et al., 2023a; Hou et al., 2022). Additionally, some research has approached the analysis and detection of under-trained tokens from a tokenization perspective (Rumbelow and Watkins, 2023; Land and Bartolo, 2024). To our knowledge, we are the first to explore token-level data selection for large language model training, aimed at enhancing data quality and information density at the most fundamental granularity.
@@ Appendix H Self-Reference Setting
where $t_{k}$ represents the i-th token in the vocabulary, and $V$ represents the size of the vocabulary. The intuition of this strategy is that the higher the information entropy, the higher the uncertainty of the token in the context. Therefore, we consider that if the language model is still uncertain for certain tokens after pretraining, we do not hope that the language model to learn it during pretraining. In Table 4, we provide more SLM results, including different select ratios and combinations of two score functions, for the convenience of readers to refer to.
@@ Appendix I Weak-to-Strong Generalization
Apart from the main experiments where we use the same base model for the reference and continual pretraining, we also investigate if a smaller reference model can effectively guide the pretraining of a larger model. We use Tinyllama-1.1B as reference model and continual pretraining Llama-2-7B on 15B OpenWebMath tokens. Results presented in Table 5 indicate that, despite the considerable gap between the small and large models (Li et al., 2023e), employing the small reference model to token selection can still yield benefits to the pre-training of the larger model. If reference and training models have different vocabularies, one can consider performing token alignment (Wan et al., 2024; Fu et al., 2023), which we leave for future work.

### v4

@@ 1 Introduction
Based on these analyses, we introduce Rho-1 models trained with a novel Selective Language Modeling (SLM) objective 111“Rho” denotes selective modeling of tokens with higher information “density ($\rho$)”.. As shown in Figure 2 (Right), this approach inputs the full sequence into the model and selectively removes the loss of undesired tokens. The detailed pipeline is depicted in Figure 4: First, SLM trains a reference language model on high-quality corpora. This model establishes utility metrics to score tokens according to the desired distribution, naturally filtering out unclean and irrelevant tokens. Second, SLM uses the reference model to score each token in a corpus using its loss (§2.2). Finally, we train a language model only on those tokens that exhibit a high excess loss between the reference and the training model, selectively learning the tokens that best benefit downstream applications (§2.2).
@@ B.2 Data Selection
Sample-level selection has been extensively studied in previous research (Sener and Savarese, 2017; Killamsetty et al., 2021), particularly through online batch selection (Loshchilov and Hutter, 2015; Schaul et al., 2015; Chang et al., 2017; Katharopoulos and Fleuret, 2018; Jiang et al., 2019). These approaches have been applied to various classification tasks (Song et al., 2020; Mindermann et al., 2022) and language modeling (Fan and Jaggi, 2023). However, Kaddour et al. (2023) find that batch selection is not computationally efficient.
Many previous works have employed the general idea of using a reference model as a proxy for data selection. For instance, Selection Via Proxy trains a proxy model to select samples with high uncertainty (Coleman et al., 2019). Xie et al. (2024a) and Engstrom et al. (2024) utilize n-gram models or datamodels with a target dataset to estimate importance weights. Additionally, Xie et al. (2024b) optimize the worst-case excess loss (Oren et al., 2019) relative to a reference model to determine domain weights. One of SLM’s scoring functions is excess loss, and the most relevant work related to excess loss is RHO-LOSS (Mindermann et al., 2022), which trains a small model on a holdout set and uses the difference between training loss and holdout loss to select in-batch samples. Although excess loss is mathematically identical to RHO-LOSS, SLM differs in three important ways: 1) The focus is distinct. Motivated by the training dynamics of token loss, the core idea of SLM is to select useful tokens for pre-training. Its score functions are highly flexible and not limited to excess loss (see Appendix H for other functions). In contrast, RHO-LOSS aims to mathematically derive a reducible holdout loss to minimize generalization loss. 2) The meaning and training procedure of the proxy model are different. SLM trains a reference model on high-quality data to reflect the desired data distribution, whereas RHO-LOSS trains a small model on a random holdout set. 3) The selection scale and granularity vary. RHO-LOSS selects sample-level data on a small scale (typically 1K–1M samples) for task-specific fine-tuning tasks such as MNIST (LeCun et al., 1998) and SST-2 (Socher et al., 2013). In contrast, SLM conducts fine-grained token-level selection on large-scale language model pre-training, involving up to 80B tokens.
Token-level training strategies have also been explored, especially for the pre-training of BERT-like models using Masked Language Modeling (MLM) (Devlin et al., 2018). Specifically, “selective masking” involves masking important tokens in the input to focus on learning tokens that are more relevant to downstream tasks (Gu et al., 2020; Lad et al., 2022), whereas “token dropping” aims to reduce training costs by omitting less important tokens (Zhong et al., 2023a; Hou et al., 2022). (Li et al., 2023e) assesses the quality of each token based on the skewness of its predicted distribution and truncates the noisy tokens during training. Additionally, some research has approached the analysis and detection of under-trained tokens from a tokenization perspective (Rumbelow and Watkins, 2023; Land and Bartolo, 2024). To our knowledge, we are the first to explore token-level data selection for large language model training, aimed at enhancing data quality and information density at the most fundamental granularity.
@@ Appendix H Self-Reference Setting
where $t_{k}$ represents the i-th token in the vocabulary, and $V$ represents the size of the vocabulary. The intuition of this strategy is that the higher the information entropy, the higher the uncertainty of the token in the context. Therefore, we consider that if the language model is still uncertain for certain tokens after pretraining, we do not expect that the language model will learn it during pretraining. In Table 4, we provide more SLM results, including different select ratios and combinations of two score functions, for the convenience of the readers to refer to.
@@ Appendix I Weak-to-Strong Generalization
Apart from the main experiments where we use the same base model for the reference and continual pretraining, we also investigate if a smaller reference model can effectively guide the pretraining of a larger model. We use Tinyllama-1.1B as reference model and continual pretraining Llama-2-7B on 15B OpenWebMath tokens. Results presented in Table 5 indicate that, despite the considerable gap between the small and large models (Li et al., 2023f), employing the small reference model to token selection can still yield benefits to the pre-training of the larger model. If reference and training models have different vocabularies, one can consider performing token alignment (Wan et al., 2024; Fu et al., 2023), which we leave for future work.